A square cost of resistance 4ohm ,50 turns and side length 5cm is places with it's plane making on amgle of 40° with a uniform field 2T.In 4.5 s the coil rotates until it's plane becomes parallel to the.magnetic field.Find the current induced in the coil.


1
Expert's answer
2022-04-11T12:03:54-0400

The current induced in the coil

"I=\frac{V}{R}=\frac{\Delta \Phi}{R\Delta t}=\frac{NBA(1-\cos\theta)}{R\Delta t}"

"I=\frac{50*2*0.05^2(1-\cos40^\circ)}{4*4.5}=3.25*10^{-3}\:\rm A"